# Do graph neural network states contain graph properties?

### v2

@@ 4 Methodology
For each of the three datasets, we use a similar network architecture consisting of a number of GNN (GCN, GIN, or GAT) layers, followed by a pooling operation (mean- (Kipf & Welling, 2017), sum-(Xu et al., 2019), or max-pooling (Hamilton et al., 2017)), and then a number of dense layers. We optimized the hyperparameters to obtain the best models possible. For the Grid-House dataset the complete hyperparameter information can be found in table 4. We ran each model 20 times and took the one with the best accuracy. As expected, there is a correlation of 0.992 between task accuracy and the maximum probing accuracy as seen in fig. 6.
@@
We compared different regularization methods. The explicit $L_{2}$ regularization encourages the network to keep the weights small, and we expect that this would make the embeddings less sensitive to fluctuations in the input data which would translate in later layers being more selective to graph properties. Dropout randomly disables a fraction of the neurons during each training iteration which forces the network to learn redundant representations, as any neuron could be dropped out. These redundant representations might make it more difficult to linearly separate the graph embeddings. We expect later layers to distinguish less between graph properties. We plot only post-pooling layers for the sake of clarity.
For the ClinTox Molecular dataset, we ranged the number of layers from 4 to 6 and hidden dimensions from 64 to 256. The final model architectures were selected based on optimal performance on the ClinTox dataset. For fMRI FC connectomes the hyperparamter search space is described in table 15.
Figure 3: Visualization of test set graph in ASD dataset with node sizes based on (from left to right): (a) degree, (b) betweenness centrality, (c) eigenvector centrality, and (d) clustering coefficient.
Probing is performed on the train and test sets, where train features $\{f_{\text{train}}^{(i)}\}$ and graph properties $\{z_{\text{train}}^{(i)}\}$ are paired for each graph (equally for the test set). Let’s define at least one example for the GCN model. Let $\mathcal{G}^{i}=(A^{i},X^{i})$ denote the $i$-th graph, where $A^{i}$ is the adjacency matrix and $X^{i}$ is the node feature matrix as previously defined. The GCN layers iteratively update the node features $H^{(l)}$ through graph convolutions defined previously as $H^{(l+1)}=\sigma(\hat{A}H^{(l)}W^{(l)})$, where $\hat{A}$ is the normalized adjacency matrix, $W^{(l)}$ are the trainable weights, and $\sigma$ is a non-linear activation function (ReLU). The node embeddings $H^{(l)}$ at each layer $l$ capture both local and global structural information by aggregating features from neighboring nodes. The final node embeddings $H^{(4)}$ are pooled using global max pooling to generate a graph-level embedding $H_{\text{global}}$, which is passed through three fully connected layers to produce the final prediction $\hat{y}$. We define these post pooling operations as $H_{\text{global}}^{(5)}=\sigma(W_{1}H_{\text{global}}),H_{\text{global}}^{(6)}=\sigma(W_{2}z_{1}),\hat{y}=\text{softmax}(W_{3}z_{2})$. For probing purposes, we use $H^{(l)}$ at different layers to evaluate node-level properties, while $H_{\text{global}},H_{\text{global}}^{(5)},H_{\text{global}}^{(6)}$, and $\hat{y}$ are used to assess graph-level properties.
We aggregate node embeddings across all graphs to train a single probing classifier for each graph property. For each property, we construct a feature matrix by combining embeddings across all graphs, layer per layer. The classifier $g$ is then trained on this aggregated dataset to predict graph properties $z_{k}^{(i)}$, where $i$ denotes the $i$-th graph and $k$ represents the $k$-th graph property, as defined in table 3. This approach assumes that the relationships between node or graph embeddings and properties are consistent across graphs.
[…]
@@
To address this, we compare two methods. One where we mean-pool the node embeddings before probing them like in fig. 12. But we might miss relevant basic properties. At the same time, we make sure that our probe is not extrapolating more structural information that what is encoded in juxtaposed node embeddings. A probing classifier, by operating on the set of embeddings simultaneously, could learn to infer relationships and patterns between nodes, leveraging the contextual relationships implicit in their embeddings. The other one where we first sort the embeddings in descending order based on their norms before concatenating fig. 4. This ordering depends only on the inherent properties of the embeddings themselves, not on their original ordering in the graph. As such, it inherently respects permutation invariance because reordering the nodes does not affect their norms or the resulting sorted order. Conveniently, sorting in this way ensures that any padding zeros align at the end of the sequence, enabling learnable representations for graphs with varying node counts. While sorting for permutation invariance is not widely discussed in the literature, it provides a practical solution by using the embeddings’ properties to enforce consistent ordering across graphs.
@@ 5.1 Grid-House dataset
The models achieved high performance scores, ranked according to their expressivity: GCN scored 0.90, GAT 0.97, and GIN 1.00 (table 5). The probing results demonstrate that the number of squares consistently yields the highest $R^{2}$ scores across all models in the global graph embeddings, both with mean-pooled GNN layer and without. This confirm our initial hypothesis where the number of square is the property of interest to perform this classification, as show the correlation between the highest R2 scores (# squares) and the performance of models fig. 6.
Figure 4: T-SNE visualisation across different layers of our GIN architecture aligned with the probing $R^{2}$ scores plots (Grid House)
When comparing the mean-pooling and norm-sorting methods, the key observation is that mean-pooling renders the probe unable to predict basic properties such as the number of nodes. It significantly reduces the probe’s ability to predict clustering and centrality-related properties and slightly diminishes its performance in predicting higher-order properties, such as the # square.
In general, the spectral and spatial perspectives are confirmed as deeper layers further smooth the signal, capturing global properties but potentially losing high-frequency details. Analogically, they extend their receptive field, it’s well visible with the message passing GCN fig. 8, incorporating more distant nodes into each node’s representation as k-hop neighborhood are taken into account. As seen in the T-SNE visualizations figs. 8, 12 and 17, the final layers tend to focus on the # squares, effectively partitioning the graphs into two classes: those with #$\#\text{squares}<5$ (indicating either the grid or house alone) and those with #$\#\text{squares}=5$ (indicating the presence of both substructures). This reduction in feature space through the layers aligns with the model’s goal of optimizing the decision boundary for binary classification, where the number of squares becomes a clear and dominant factor for separability. Interestingly enough, the GIN model has its layer MLP1 partitioning the graphs label 0 between those with a house and those with a grid using # of triangle significantly. The Gat model present a form of confusion between classes at the layer Global before clearly sharping its representation and classification abilities. The other present properties, density and average path length are also prominent and make sense as the presence of both a house and a grid does slightly increase the average density and path length of graphs. These findings confirm the correspondence between graph embeddings clustering and property hyperplane separation.
We further observe that, for both the GCN and GIN models, the application of $L_{2}$ regularization yields the expected behavior. The last layer of the GCN in fig. 10 shows a stronger dominance of the number of squares feature when $L_{2}$ regularization is applied compared to when it is not. Similarly, in the GIN fig. 14, both # triangles and density become less detectable relative to the number of squares, by the probing classifier in the final layers under $L_{2}$ regularization, consistent with the anticipated effects on the feature representation. Dropout leads to the representation of multiple properties in the final layers. Notably, in fig. 11, the last layer exhibits a reduced separability gap between the # square and other properties, indicating a more distributed feature representation when dropout is applied.
We can use the number of square R² score of models to compare their expressivity. For the GCN fig. 8, the slow increase of the R² score correspond to the message-passing theory as we said. The use of the density as a side property make sense but also indicate the lack of optimization of this model, which we can attest by looking at the T-SNE of the last layer where some graphs are misgrouped. The regularization seems to correct this effect partialy fig. 10.
The GAT and GIN models shows a strong dominance of # of square from the beginning. Either with an Attention-Weighted Aggregation or a MLP injective aggregation, we don’t expect such score from the first layer. The receptive field is limited to 1-hop nodes. This could be explained by a form of signal short-cut where local features like the degree of neighbors is already enough to expect a four-cycle pattern. Similarly, the closeness centrality of a node or its clustering coefficient is barely encoded in the first layers table 7. The effectivity of both the GAT and the GIN models would then catch these short-cuts very effectively.
[…]
@@ 5.2 ClinTox Molecular
As expected, the GIN model outperform the other models with a test accuracy of 0.93 table 8. When looking at the corresponding linear probing performance in table 1, we find that the highest scores are consistently yielded by the average degree, the spectral radius, the algebraic connectivity and the density, in that order.
@@
GIN Layer Avg. degree Spectral radius Alg. co. Density Avg. btw. cent. Graph energy
@@
These findings validate our methodology on already known domain knowledge. Indeed, the average degree of atoms in a molecule provides a straightforward interpretation, as atoms with higher valencies are generally less stable and less biologically compatible. For instance, hydrogen with a valency of 1 and oxygen with a valency of 2 are more compatible with carbon-based molecules, whereas sulfur, with a valency of 6, is less favorable for biological systems  (Komarnisky et al., 2003). Therefore, the average degree serves as a useful indicator of molecular toxicity. Additionally, the spectral radius, often associated with molecular stability and reactivity, is another valuable graph property. Molecules with a lower spectral radius tend to be more stable, while those with a higher spectral radius may exhibit localized electron densities, increasing their reactivity. Using this property to predict molecular toxicity is a logical approach.
5.3 fMRI FC connectomes
Here, our primary focus is to uncover properties under exploited by the field. GIN outperforms other architectures table 14, reproducing the observation by Zheng et al. (2023). It makes sense to study exclusively the last layers of the GIN model. The probing results on the ASD dataset reveal that the number of triangles consistently achieves high $R^{2}$ scores across all models, with particularly strong performance in GIN models. This property is followed by the spectral radius and the density.
For the MDD results, detailed results tables 20, 23, 24 and 22 reveal that the number of triangles still consistently achieves high $R^{2}$ scores across all models while being less of a distinctive feature than in ASD. This time, the spectral radius is dominated by the density of the graph. In general, the embeddings from the 7th layer of our GIN architecture exhibit higher $R^{2}$ scores for relevant graph properties, suggesting improved separability in the embedding space for MDD classification compared to ASD. This indicates that the learned representations at this depth capture more discriminative structural features, facilitating more effective class separation between MDD and healthy controls.
@@ 6 Discussion
6.1 Expectations
With Grid-House we hypothesized that the GNNs would benefit from leveraging the # of square to render the problem linearly separable. Based on their mathematical restrictions, we hypothesized that the GIN would perform better than the GAT and the GCN. Regularization methods should either refine the representations, as seen with $L_{2}$ regularization, or distribute them more broadly, as achieved with dropout.
@@ 6.1 Expectations
Based on the message-passing paradigm, we anticipated a clear absence of the # of square in the first layer. Additionally, we expected that the mean-pooling and norm-sorting methods would not significantly alter how representations are probed, except for basic properties like the number of nodes, which are easily interpretable from the tensor of node vectors but not from an aggregated representation.
For the ClinTox Molecular dataset, based on the literature (Kengkanna & Ohue, 2024; Chen et al., 2021; Jiang et al., 2021) some few properties have been found to be link with toxicity such as the node degree (i.e. the valency), subgraph patterns (functional groups, chemical fragments), and the overall graph connectivity.
Based on existing literature on functional connectivity (FC) network properties in ASD and MDD appendix G, we hypothesized that specific properties would be critical in classifying brain networks for the fMRI FC connectomes dataset. For ASD, we expect betweenness centrality to play a significant role at the node level, reflecting local overconnectivity. At the graph level, we anticipate that clustering coefficient, characteristic path length, and small-worldness will be essential in capturing the local and global network disruptions seen in ASD, particularly the imbalance between local overconnectivity and long-range underconnectivity. For MDD, we hypothesise that increased clustering coefficients, modularity, number of triangles and number of squares will be key features for classification, as they could indicate of heightened local interconnectedness and disrupted global integration.
6.2 Findings
@@
We first demonstrate the feasibility of our probing method through the Grid-House dataset. In line with our expectations, The number of squares metric dominated across all layers and models, with GIN showing enhanced expressivity. Secondary properties like avg_path_lenght figs. 7 and 8 showed early significance but gradually diminished through the layers, demonstrating how GNNs act as low-pass filters on graph signals, the receptive field select the most discriminative properties. These findings align with the labels, as the density or the average path length is increase when both a grid and a house are added to a graph. Mean-pooling and norm-sorting methods exhibit different behaviors that can be explained by the type of information the classifier access : aggregated graph data or detailed node embeddings. The probe might also infer relationships and patterns that are implicitly encoded in individual node embeddings but become explicit when considered collectively.
The analogy with computer vision lead to the question of supervision signal as we would expect unsupervised model to capture structural information more effectively than supervised ones (Kingma & Welling, 2022). Unsupervised learning algorithms can discover hidden patterns and structures in data without explicit instructions, potentially uncovering relationships that may not be immediately obvious to humans. Following this logic, this ability to find latent structures could be advantageous in capturing complex structural information in graph data and lead to similar breakthought than discovering new particle physics phenomena using clustering and dimensionality reduction on collider data or identifying subtypes of cancer with clustering and PCA in genomic studies.
@@
With the fMRI FC connectomes dataset, the results provided new insights that extended beyond our initial hypothesis. The prominence of the number of triangles highlighted the importance of local structural motifs. This makes sense in the context of functional connectivity, where local overconnectivity in specific brain regions, such as sensory and association cortices, has been observed in individuals with ASD. The strong role of triangle motifs may reflect the tight, redundant local connections that characterize these regions, supporting the hypothesis that local overconnectivity is a key factor in ASD. The spectral radius and density and graph energy being particularly significant is also logical, as these properties are closely related to the overall connectivity strength and the compactness of connections within subnetworks. The outcomes for both the ASD and the MDD datasets showed promising results that should be discussed more deeply with neuroscientists. The results mainly suggest the importance of graph substructures or spectral and small-world properties over more basic graph properties to explain how graph neural networks predict these neurological disorders in the FC matrices of patients. This is a result which, to the best of our knowledge, has never been investigated before.
6.3 Future work
Our methodology has several limitations. While we addressed dataset issues such as leakage and isomorphic graphs, a key challenge remains the lack of guarantees that GNNs find globally optimal solutions, despite their theoretical capacity as universal function approximators. This is particularly evident in fMRI data, where multiple layers of complexity—from MRI limitations and BOLD signal characteristics to Pearson correlation for functional connectivity—introduce noise and inaccuracies. Investigating additional graph properties like girth or complex motifs could be beneficial. Preliminary work on alternative architectures (e.g., GATv2, GraphSAGE, ChebNet, Set2Set, HO-Conv, DiffPool) has begun but is not yet complete. An extensive exploration of 1-WL, 2-WL and 3-WL GNN equivalent could bolster the paper’s contributions by showing clear restrictions and capabilities of these models. As a future encouraging work, the question of supervision signal would be very insightful in understanding what type of properties are learn by unsupervised models.
@@ 7 Conclusion
We demonstrate the relevance of using probing classifier as a model-agnostic explainability method for graph neural networks. We manifest both the expressivity of different GNN architectures and their ability to solve a graph classification problem through optimal feature extraction. They render it linearly separable in the space of their embeddings through the computation of graph properties. We validate domain knownledge with the Clintox Molecular dataset and investigate the possibility of formulating hypotheses on the emergent dependence of complex systems attributes to basic and more higher level structural properties with the fMRI FC connectomes dataset. We explore how the macroscopic behavior of complex systems emerges from the intricate interplay of their microscopic components. There is a manifest emergence of molecular qualities like toxicity with regard to their structural properties like node degree (atom valency) and spectral radius (the molecule’s stability). To explain a macro attribute, there are instances where structural properties may offer more insight than the mere aggregation of element properties. Both can be interconnected, yet the challenge remains in disentangling these relationships.
@@
Pelletreau-Duris Tom: Conceptualization (lead); methodology (lead); software (lead); formal analysis (lead); writing – original draft (lead); writing – review and editing (equal). Ruud van Bakel: writing – review and editing (equal). Michael Cochez:Writing – review and editing (equal).
Code
All results and experiments accessible on github
@@ Appendix A R2 Score
[…]
@@
[…]

### v3

@@ 4 Methodology
We propose a pipeline to confirm that graph-theoretic properties are a good candidate for studying GNN inductive bias. For each of the three datasets, we compare different GNN architectures (GCN, GIN, or GAT). The first layers are followed by a pooling operation (mean- (Kipf and Welling, 2017), sum-(Xu et al., 2019), or max-pooling (Hamilton et al., 2017)), and then a number of dense layers. For the Grid-House dataset the full hyperparameter settings after optimization can be found in table 4. Because we can only probe on one model weights, we ran each model 20 times and took the one with the best accuracy. There is a correlation of 0.992 between the accuracy and the maximum probing score fig. 6.
@@
We also compare different regularization methods and their effect on GNN representations. It is known that explicit $L_{2}$ regularization encourages the network to keep the weights small. We expect that this will make the embeddings less sensitive to fluctuations in the input data which would translate in later layers being more selective to graph properties, leading to more specific receptive fields. Dropout randomly disables a fraction of the neurons during each training iteration which forces the network to learn redundant representations, as any neuron could be dropped out. These redundant representations might make it more difficult to linearly separate the graph embeddings. We expect later layers to distinguish less between graph properties. In other words, we expect more polysemanticity (Bereska and Gavves, 2024). We plot only post-pooling layers for the sake of clarity.
As a second set of experiments, we use our build pipeline to probe GNNs on real-life datasets, wondering if what we find aligns with domain knowledge in chemistry. For the ClinTox Molecular dataset, we ranged the number of layers from 4 to 6 and hidden dimensions from 64 to 256. The final model architectures were selected based on optimal performance on this dataset. We also did preliminary work on Knowledge Graphs and fMRI connectomes, which is available in the appendix.
@@
Before probing, we first sort the embeddings in descending order based on their norms before concatenating fig. 3. This ordering depends only on the inherent properties of the embeddings themselves, not on their original ordering in the graph. As such, it inherently respects permutation invariance because reordering the nodes does not affect their norms or the resulting sorted order. Conveniently, sorting in this way ensures that any padding zeros align at the end of the sequence, enabling learnable representations for graphs with varying node counts.
@@ 5.1 Grid-House dataset
All tested models achieve a test accuracy of over 0.9, with the GIN without regularization or dropout reaching the highest score of 1.0 (see table 5 in the appendix for more details). The probing results in fig. 3 demonstrate that the number of squares consistently yields the highest $R^{2}$ scores. This focus on #squares, effectively partitions the graphs into two classes: those with $\text{squares}<5$ (indicating either the grid or house alone) and those with $\text{squares}=5$ (indicating the presence of both substructures). We see the same results for the GAT model (fig. 4), especially for the final layers and for other model variants (see the appendix). This confirms our initial hypothesis where the number of square is the property of interest to perform this classification. Interestingly, the MLP1 layer of the GIN model can effectively separate the graph embeddings using the #triangles feature.
Figure 3: GIN probing results $R^{2}$ across different layers aligned with the T-SNE visualisations of the embedding (Grid House)
@@ 5.2 ClinTox Molecular
As expected, the GIN model outperform the other models with a test accuracy of 0.93 (table 8 in the appendix). We find that the highest performance is consistently achieved by probes for the features of average degree, the spectral radius, the algebraic connectivity and the density, in that order (see table 1).
These findings validate our methodology on already known domain knowledge. Indeed, the average degree of atoms in a molecule provides a straightforward interpretation, as atoms with higher valencies are generally less stable and less biologically compatible. For instance, hydrogen with a valency of 1 and oxygen with a valency of 2 are more compatible with carbon-based molecules, whereas sulfur, with a valency of 6, is less favorable for biological systems  (Komarnisky et al., 2003). Therefore, the average degree serves as a useful indicator of molecular toxicity. Additionally, the spectral radius, often associated with molecular stability and reactivity, is another valuable graph property. Molecules with a lower spectral radius tend to be more stable, while those with a higher spectral radius may exhibit localized electron densities, increasing their reactivity. Domain knowledge on molecular toxicity align with the GIN strategy.
5.3 Pooling
While have deferred the results for regularization and dropout to the appendix, we did notice that regularization tends to help separate out the #squares better, while dropout has the opposite effect. This is in line with our expectations.
@@ 5.3 Pooling
GIN Layer Avg. degree Sp. radius Alg. co. Density Avg. btw. cent. Graph energy
@@ 6 Discussion
Expectations: With Grid-House we hypothesized that the GNNs would benefit from leveraging the # of square to render the problem linearly separable. Based on their mathematical restrictions, we hypothesized that the GIN would perform better than the GAT and the GCN. Regularization methods should either refine the representations, as seen with $L_{2}$ regularization, or distribute them more broadly, as achieved with dropout. Based on the message-passing paradigm, we anticipated a clear absence of the # of square in the first layer. Additionally, we expected that the mean-pooling and norm-sorting methods would not significantly alter how representations are probed, except for basic properties like the number of nodes, which are easily interpretable from the tensor of node vectors but not from an aggregated representation. For the ClinTox Molecular dataset, based on the literature (Chen et al., 2021; Jiang et al., 2021; Kengkanna and Ohue, 2024) some properties have been found to be link with toxicity such as the node degree (i.e. the valency), subgraph patterns (functional groups, chemical fragments), and the overall graph connectivity.
@@
Findings: We first demonstrate the feasibility of our probing method through the Grid-House dataset. In line with our expectations, The number of squares metric dominated across all layers and models, with GIN showing enhanced expressivity. Secondary properties like avg_path_lenght figs. 7 and 8 showed early significance but gradually diminished through the layers, demonstrating how GNNs act as low-pass filters on graph signals. The receptive field selects the most discriminative properties. These findings align with the labels; density or average path length increases when both a grid and a house are added to a graph.
The inductive bias of Graph Neural Networks (GNNs), when viewed through the lens of the linear representation hypothesis, reflects a tendency to organize internal representations such that task-relevant features become increasingly linearly decodable as a function of the supervision signal. In this framework, the supervision signal guides the layer-wise development of representations, selectively reinforcing directions in activation space that align with informative graph-theoretic properties (e.g., centrality, motif counts, degree distributions).
@@
7 Future work
Our methodology has several limitations. While we addressed dataset issues such as leakage and isomorphic graphs, a key challenge remains the lack of guarantees that GNNs find globally optimal solutions, despite their theoretical capacity as universal function approximators (Hornik et al., 1989). As we observed, early layers in the network contain predictive information for some global graph features. It is however unclear why this is the case. They might be somehow predictive for the features we are looking for, perhaps in the specific combination with the specific distribution of graphs we used int he experiments. Investigating additional graph properties like girth or complex motifs could be beneficial. Preliminary work on alternative architectures (e.g., GATv2, GraphSAGE, ChebNet, Set2Set, HO-Conv, DiffPool) has begun but is not yet complete. An extensive exploration of 1-WL, 2-WL and 3-WL GNN equivalent could bolster the paper’s contributions by showing clear restrictions and capabilities of these models. As a future encouraging work, studying the supervision signal comparing unsupervised, self-supervised and supervised models would be very insightful. Interesting application research could be done on Knowledge Graphs. Another path could be fMRI connectomes, which are also graphs. With fMRI data, some functional connectivity patterns could be associated with cognitive states under the lens of GNNs probing.
@@ 8 Conclusion
We demonstrate the relevance of using a probing classifier as a model-agnostic explainability method for graph neural networks. We manifest both the expressivity of different GNN architectures and their capacity to solve a graph classification problem through effective feature extraction. They render it linearly separable in the space of their embeddings through the computation of graph properties. We validate domain knowledge with the Clintox Molecular dataset. There is a manifest emergence of molecular qualities like toxicity with regard to their structural properties like node degree (atom valency) and spectral radius (the molecule’s stability). To explain a macro attribute, there are instances where structural properties may offer more insight than the mere aggregation of element properties. This provide encouraging results for studying the possibility of formulating hypotheses on the emergent dependence of complex systems attributes to basic and more higher level structural properties. We could explore how the macroscopic behavior of complex systems emerges from the intricate interplay of their microscopic components, raising hypothesis on graph similarly than what is now done on language with LLMs or on vision with computer vision.
Acknowledgment
This work is based on the MSc. AI thesis by Tom Pelletreau-Duris, a large part of the results were also presented in that work.
Michael Cochez is partially funded by the Elsevier Discovery Lab, partially funded by the Graph-Massivizer project, funded by the Horizon Europe programme of the European Union (grant 101093202), and supported by a gift from Accenture LLP. His work on this publication is in part based upon work from COST Action CA23147 GOBLIN - Global Network on Large-Scale, Cross-domain and Multilingual Open Knowledge Graphs, supported by COST (European Cooperation in Science and Technology, https://www.cost.eu).
We would also want to thank the anonymous reviewers for their comments and suggestions that helped us improve the manuscript.
@@ Acknowledgment
Pelletreau-Duris Tom: Conceptualization (lead); methodology (lead); software (lead); formal analysis (lead); writing – original draft (lead); writing – review and editing (equal). Ruud van Bakel: writing – review and editing (equal); Michael Cochez: writing – review and editing (equal); Supervision; Funding acquisition
@@ Appendix A R2 Score
We are using $R\textsuperscript{2}$ as the main metrics. The $R\textsuperscript{2}$ score (coefficient of determination) measures the proportion of variance in the dependent variable that is predictable from the independent variable(s). For a probing classifier, it would indicate how well the probe’s predictions match the actual properties being probed. More formally, $R\textsuperscript{2}$ is defined as:
@@
An $R\textsuperscript{2}$ value ranges from 0 to 1, where: $R\textsuperscript{2}=1$ means the probing classifier perfectly explains the variance in the target property (i.e., the learned representations fully capture the property). $R\textsuperscript{2}=0$ means the probing classifier does no better than predicting the mean $\bar{z}$, implying the representations do not capture any useful information about the property. Good R2 score should indicate how the model achieves its behavior on the original task Hupkes et al. (2018).
A good $R\textsuperscript{2}$ score gives a sense of how well the features at each layer can be separated linearly to predict the target labels. The second reason is that a more complex probe “bears the risk that the classifier infers features that are not actually used by the network” (Hupkes et al., 2018). Of course, other non linear probes have been explored in the literature (Belinkov, 2021). If a few studies observed better performance with more complex probes, the logic remained the same $\operatorname{Perf}\left(g,f_{1},\mathcal{D}_{O},mathcal{D}_{P}\right)>\operatorname{Perf}\left(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P}\right)$, of two representations $f_{1}(x)$ and $f_{2}(x)$, holds across different probes $g$. The important criteria is to compare the results obtained by the same measurement system. In general, if we can predict one property on one embedding for a given classification problem, then it means this properly is useful for the problem resolution.